OpenStudy (anonymous):

What is the slope and y-intercept of the line represented by this equation? -3x+8y=-24

OpenStudy (istim):

IT's right there in the equation.

OpenStudy (anonymous):

It probably is, but I suck at math. So I don't know the slope or y-intercept of this equation.

OpenStudy (istim):

Ok. In the formula y=mx+b, m is the slope, and b is the y-intercept. This formula is just a variation of that formula.

OpenStudy (anonymous):

so would the slope be -3/8 and would the y-intercept be -8?

OpenStudy (istim):

Eh, I don't know. I personally didn't do the work. I guess it works, if you rearranged -3x+8y=-24 to become y=mx+b (gotta remove the 8 on y remember). If you did that, then you are correct.

OpenStudy (istim):

You good then?